162 Preliminary Report

Site 987

While underway to Site 987 (EGM-4), we received conflicting ice map data regarding the outer edge of the ice margin. The most recent SSMI data indicated that the ice edge was well to the west of the proposed EGM-4 site. Later, advanced very-high-resolution radiometer (AVHRR) and SAR data showed the site to be 10 nmi within closed ice. As a result, the first survey waypoint was selected well to the east along the 70°30.0' latitude. At 2400 hr on 22 August 1995 the vessel slowed to survey speed (5.4 kt). In dense fog a seismic profile toward the ice margin in the west was begun. Repeated freezing of the water guns hampered the survey and eventually all water guns were inoperative. The profile extended all the way to 18°9.89'W. This data allowed a tie with an existing pre-cruise survey line that crossed the proposed EGM-4 site. The location of the ice margin was estimated to be approximately 10 nmi farther to the west, or roughly at the coordinates of the original EGM-4 drill site. The vessel used GPS coordinates to return to a previously approved alternate site, EGM-4B, located on the site survey line. Unfortunately, a cross line could not be shot over site EGM-4b due to the gun failures.

The operational plan for this site was similar to that for Site 986 (SVAL-1B). It called for two piston-cored holes drilled to refusal or approximately 200 m below the sea bed. One hole was to be deepened to XCB refusal, which was estimated to occur at about 500 mbsf. A third hole was to be drilled down just short of the XCB hole TD, and then continuously RCB-cored to a depth of 800 mbsf. This hole would terminate shy of basement, but penetrate the glacial sediment package which was the major site objective. A minimum number of downhole temperature measurements were to be taken in the upper (APC) portion of the hole in order to establish a temperature gradient. A full suite of logging tools was to be deployed on the deepest site. Iceberg warning and hole abandonment procedures were put in effect. In addition, the Danish Greenland Command was advised of the vessel's arrival onsite and its anticipated departure from the area.

The positioning beacon was deployed at 0920 hr on 23 August 95 initiating Hole 987A. Routine piston coring continued through Core 987A-10H. No core was recovered from Core 987A-3H, because the core-catcher flapper stuck open, allowing the entire core to slip out of the liner. Since APC refusal occurred at such a shallow depth, coring operations were continued with the XCB-system. XCB coring continued with excellent recovery until coring was halted after Core 987A-22X, when the scientific target for the hole was reached. Methane was encountered ranging from 7 to 67760 ppm and ethane from 0 to 7 ppm. Propane was detected beginning with Core 987A-13H and ranged from 0 to 1 ppm. Higher molecular weight hydrocarbons were not detected.

After offsetting the vessel 15 m east, a bottom-water temperature was taken with the APC temperature tool before spudding Hole 987B. APC coring continued through Core 987B-11H when the scientific target depth for the hole was reached. Core recovery was lost for Cores 987B-3H through 5H, again because of the core-catcher flapper sticking open. Four successful downhole temperature measurements were taken on Cores 987B-4H, 6H, 8H, and 10H, yielding a temperature gradient of 96°C/km.

Because of the lost core interval on the first two holes, a third APC hole, Hole 987C, was drilled through the problem zone. This time the core-catcher flapper was supplemented with a dog-type core catcher as well. Running dual core catchers worked successfully, and coring was suspended once the target for the hole was reached.

The ship was again offset 15 m east. Routine piston coring continued without incident at Hole 987D through Core 987D-11H. XCB coring was initiated with Core 987D-12X. The hard formation XCB cutting shoe was severely damaged on Core 987D-21X, apparently due to coring through a large dropstone. No core was recovered through an apparent mass-flow deposit from Cores 987D-21X through 23X. Normal coring parameters were regained on Core 987D-24X. Coring was ultimately discontinued due to a steadily deteriorating rate of penetration (ROP) and core recovery. In addition, the cores became increasingly more disturbed and many core "biscuits" were noticed. Cores 987D-24X through 42X contained methane from 3162 to 21087 ppm and ethane from 2 to14 ppm. Propane ranged from 0 to 3 ppm. Higher molecular weight hydrocarbons were not detected.

The first indications of encroaching sea ice were identified on the radar during the coring of Hole 987D. The floe was picked up at a range of approximately 6.3 nmi from the drill site. It was moving toward the site at approximately 0.25 kt; however, the closest point of approach was 5.5 miles, so the ice never posed a danger to drilling operations. So the drill pipe was tripped back to the drill floor to change the bit/BHA to one used for RCB coring in preparation for Hole 987E. During the pipe trip, the vessel was offset 50 m to the east. Several hundred meters of the upper formation were to be drilled rather than cored; therefore, the seafloor depth from Hole 987D was used as a reference. This is more accurate than "feeling" for bottom with the RCB drilling assembly. Drilling ahead continued to a depth of 2047.5 mbrf (363.3 mbsf). The center bit was then recovered via wireline, and continuous RCB coring commenced. Coring was terminated after Core 987E-52R at 859.4 mbsf. This depth was in excess of the original scientific target depth of 800 mbsf; however, permission was obtained earlier to advance to 900 mbsf. Coring had to be abandoned only a few tens of meters short of the elusive basement so the logging program would not be compromised. Maximum methane and ethane identified were 70,787 ppm at a depth of 792.9 m, and ethane was 62 ppm at a depth of 739.0 m.

During the following wiper trip, a tight spot was identified at 408.8 mbsf that required 30,000 lb of overpull. Then, while running back in the hole, a hard spot was tagged with 30,000 lb of bit weight at 409.8 mbsf. At that point, the top drive was picked up; however, advancement still could not be achieved so a center bit was dropped. Washing/reaming with the top drive continued from that point until the original TD for the hole was reached. An additional 60-barrel mud sweep was pumped during the reaming process. Twenty-two meters of fill were found on the bottom of the hole. Prior to recovering the center bit, the hole was swept a final time with 50 barrels of gel mud. Two additional wireline runs were then made to release the bit and reverse shift the MBR sleeve. The pipe was placed at a logging depth of only 2122.7 m (438.5 mbsf), because there was little confidence that the tools would be able to traverse the bad spot in the hole that was identified during the wiper trip. The logging equipment was rigged, and the first run, with the Quad combo suite of tools, was deployed to a depth of 2165.0 m (480.8 mbsf). This was only 42.3 m below the open-ended pipe, so the logging tools were removed and the pipe was lowered to a depth of 2180.5 m (496.3 mbsf). The second run with the Quad combo tools reached a depth of 2184.0 m (499.8 mbsf), which was only 3.5 m below the open-ended pipe. Further logging attempts in the lower part of the hole were considered futile. The tools were recovered, the logging sheaves were again partially rigged down, and the pipe was pulled to a depth of 1776.1 m (91.9 mbsf). A third run in the upper portion of the hole with the Quad combo suite of logging tools reached a depth of 2170.0 m (485.8 mbsf) or 394.0 m below open-ended pipe. This logging run was successfully completed at 1935 hr on 31 August 95.

The next logging run with the GHMT tool reached a depth of 2164.0 m (479.8 mbsf) or 387.9 m below open-ended pipe. This logging run was successfully completed at 2320 hr on 31 August 95. The final logging attempt with the FMS reached a depth of 2092.0 m (407.8 mbsf) or 315.9 m below open-ended pipe. During this final run, another set of data points documenting the performance of the new wireline compensator control system was collected for the Lamont-Doherty Earth Observatory, Borehole Research Group (LDEO/BRG). The tools were subsequently recovered on deck at 0400 hr and the logging equipment was rigged down.

The drill pipe knobbies were then laid out and the drill string was pulled clear of the seafloor at 0536 hr. The beacon was released and recovered during the subsequent pipe trip. The BHA components were broken down, cleaned, and stowed for transit. Thrusters and hydrophones were pulled and the remaining rig floor equipment was secured for the 1.7 day (455 nmi) transit to Reykjavik, Iceland. As the vessel got underway in moderate snow flurries at half speed, scattered bergy bits and a few isolated growlers, likely discharged from Scoresby Sund, littered the area. The ice margin was clearly visible to the west as the vessel sailed on a southerly course toward Iceland.

Table 1- Hole Summaries

162 Table of Contents

Publications Home

ODP Home